OpenStudy (anonymous):

write down the first three terms of the expansion (x^2 - 1/x)^-2

OpenStudy (anonymous):

power series expansion? we can do this

OpenStudy (anonymous):

is there a snap way you think or do you just grind it out? first we should write \[\frac{1}{(x^2-\frac{1}{x})^2}=\frac{x^2}{(x^3-1)^2}\]so it makes sense at 0